OpenStudy (anonymous):

medals rewarded! What is the probability of drawing four aces in a row from a shuffled standard deck of playing cards?

OpenStudy (anonymous):

not very likely \[\frac{4}{52}\times \frac{3}{51}\times \frac{2}{50}\times \frac{1}{49}\]

OpenStudy (anonymous):

hope the method is clear